OpenStudy (anonymous):

NEED hELP 5+4*2+6-2*2-1

OpenStudy (anonymous):

Im not good at this stuff

OpenStudy (derpisthewordorisit):

14

OpenStudy (anonymous):

is that 4 to the power of 2?

OpenStudy (derpisthewordorisit):

33

OpenStudy (derpisthewordorisit):

ITS 33

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

22

OpenStudy (derpisthewordorisit):

33 dude, its not 22

OpenStudy (anonymous):

4*2 is 16 and 2*2 is 4

OpenStudy (anonymous):

now solve from there

OpenStudy (derpisthewordorisit):

Do it on a caculater.

OpenStudy (derpisthewordorisit):

If you know how to use a caculater

OpenStudy (derpisthewordorisit):

I got 33,

OpenStudy (anonymous):

5+16=21+6=27-4=23-1=22 so 22

OpenStudy (derpisthewordorisit):

Look,

OpenStudy (derpisthewordorisit):

5+4*2=?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Insert parentheses so that it has a value of 19.

OpenStudy (anonymous):

so 21

OpenStudy (derpisthewordorisit):

What the hell, that's 18..

OpenStudy (derpisthewordorisit):

9x2=18

OpenStudy (anonymous):

how are you getting 9?

OpenStudy (derpisthewordorisit):

from 5+4...

OpenStudy (anonymous):

ORDER OF OPERATIONS MAKES YOU DO 4*2 FIRST

OpenStudy (derpisthewordorisit):

OH ORDER OF OPERATIONS LOL MB

OpenStudy (derpisthewordorisit):

DIDNT READ IT XD

OpenStudy (anonymous):

Parenthesis Exponents <----- Multiplication Division Addition <--------- Subtraction XD

OpenStudy (anonymous):

SO 22

OpenStudy (derpisthewordorisit):

Yeah, I know :P

OpenStudy (derpisthewordorisit):

He's right, I did the math..

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

your welcome